I have two questions so thanks to anyone who answers them
Plz provide steps as well as I have been stuck on these problems for quite a while
A garden has a perimeter of 72 ft and an area of 360. What are the length and width?
If Home Depot sells 2x4 for $15.25 a yard how much will it cost to design a border for the walkway which has a length of 45 feet and a width of 1 foot

Answer by Alan3354(69443)

A garden has a perimeter of 72 ft and an area of 360. What are the length and width?
----
P = 2L + 2W = 72
L + W = 36
L*W = 360
------------
If Home Depot sells 2x4 for $15.25 a yard how much will it cost to design a border for the walkway which has a length of 45 feet and a width of 1 foot
Is it a 2 by 4 in inches? or feet?